OpenStudy (anonymous):

After 3 days a sample of radon-222 has decayed to 58% of its original amount. (a) What is the half-life of radon-222? (Round your answer to two decimal places.)

OpenStudy (anonymous):

please understand that whenever half-life is given or asked for, the simplest equation to use is f = 0.5^(t/h) , where f = fraction remaining , and h = half-life ========= qa t/h = ln f / ln 0.5 h = t(ln 0.5 / ln 0.58) plugging in, h = 3(ln 0.5 / ln 0.58) = 3.82 days <------- qb simlarly, t = h(ln f / ln 0.5) plugging in, t = 3.82(ln 0.1 / ln 0.5) = 12.68 days <-------

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your Welcome :)

OpenStudy (anonymous):

theres actually another part to it b) How long will it take the sample to decay to 10% of its original amount? (Round your answer to two decimal places.)

OpenStudy (anonymous):

Well in general, for this question, the fomula for radioactive decay is: \[m_f=m_i\left(\frac{1}{2}\right)^{\frac{t}{H}}\] Where: \[\eqalign{ &(m_f)\textit{ represents the final mass} \\ &(m_i)\textit{ represents the initial mass} \\ &(H)\textit{ represents the half life (time required for substance to halve in mass} \\ &(t)\textit{ represents the time taken since beginning} }\] So then we don't know our initial or final amount except that \(m_f=58\%\textit{ of }m_i=0.58m_i\) We know our time is 3 days, which for practical reasons we will write is 36 hours So then we can write: \[0.58m_i=m_i\left(\frac{1}{2}\right)^\frac{36}{H}\] We can divide by \(m_i\) to obtain: \[\eqalign{ 0.58&=\left(\frac{1}{2}\right)^\frac{36}{H} \\ \log_{10}(0.58)&=\frac{36}{H}\log_{10}\frac{1}{2} \\ \frac{\log(0.58)}{\log(0.5)}&=\frac{36}{H} \\ \frac{\log(0.58)}{36\log(0.5)}&=\frac{1}{H} \\ \frac{36\log(0.5)}{\log(0.58)}&=H }\]

OpenStudy (anonymous):

Whoops lol im a little slow there...ill do the second one!

OpenStudy (anonymous):

By the way, I get 45.80 hours...

OpenStudy (anonymous):

like for A I get 45.80 hours...any ideas @sana99 ?

OpenStudy (anonymous):

And i'll do B right now @CrazyCarrie6655 :)

OpenStudy (anonymous):

Assuming that the half life is 45.8 hours, then we have: \[0.1=\left(\frac{1}{2}\right)^{\frac{t}{45.8}}\] So we isolate for t and obtain: \[45.8*\frac{\log0.1}{\log0.5}=t\approx152.14\] So therefore, it will take 152.14 hours to decay to 10% it's original amount.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No prob :)